# Mean-Shift Distillation for Diffusion Mode Seeking

### v1

@@ 4.3 Practical Setting
For large-scale image datasets, idealized denoiser is no longer tractable and we have to contend with a learned denoising function, as well as all the associated machinery.
This introduces numerical issues. Namely, as already mentioned in Section 3.4, the magnitude of the kernel term may grow to where the standard first or second order integrators can no longer manage it; but conversely, so does the magnitude of the learned score term when $z_{t}$ is far out of distribution; why this is the case becomes apparent when we consider that the ideal denoiser (Section 4.1) uses the same equation as mean shift. This is a problem because at the start of the optimization, it is likely in a high-dimensional space that $x$ will be out of distribution and we have to choose between the integration failing because the denoiser term has a high magnitude, or because the kernel term has a high magnitude.
@@ 4.4 Pre-trained Stable Diffusion
We use the latent-space diffusion model, Stable Diffusion, as the diffusion prior for text-conditioned optimization of parameters of differentiable image generators. Specifically, we optimize parameters $\vartheta$ of generator $g$, a rendering function that maps $\vartheta$ to an image $\mathcal{I}$. The rendered image $\mathcal{I}$ is fed to the image encoder to get $x^{k}$, our latent at optimization step $k$, over which the gradient is computed. We define two settings, (1) where $\vartheta$ represents an RGB image, and (2) where $\vartheta$ represents a 3D volume. Specifically:
@@
We implement all our code in PyTorch, on a single NVIDIA A100 gpu. We use the Threestudio (Guo et al., 2023) framework for experiments involving pre-trained Stable Diffusion. All the experiments use Adam optimizer with lr$=10^{-2}$. We set optimization steps to $400$ for text-to-2D and $7k$ for text-to-3D. We use a monotonically decreasing schedule for the bandwidth $\lambda$.
@@
Figure 4: FID vs optimization iterations.
@@ 4.5 Evaluation
We use a subset of the prompts curated by Poole et al. (2022); Hertz et al. (2023). We include all prompts in Appendix B.
@@
For toy density dataset (Section 4.2), we compute negative log-likelihood scores (NLL), generative precision and recall (Kynkäänniemi et al., 2019), and maximum mean discrepancy (MMD). For text-to-2D, we use images produced by DDIM to represent the ground truth distribution. To evaluate fidelity of the images, FID (Heusel et al., 2017) is computed for each baseline (SDS, SDI) and ours against this ground truth image set. We also compute CLIP scores (cli, ) to measure prompt-generation alignment.
@@
Table 3 reports results for FID and CLIP-based similarity, comparing our method with SDS and SDI. We outperform both baselines in image fidelity and are our generations are more faithful to the input prompt.
@@
Figure 2 (top row) and Figure 8 compares our method with SDS and SDI on text-to-2D generation, qualitatively. In the latter, we show the importance of the two heuristics (Section 4.3) to resolve numerical instabilities, absence of which can result in visual artifacts. SDS, as discussed, produces low-fidelity results while SDI’s inversion accumulates numerical errors during early stages of optimization. In Figure 5, we qualitatively compare results for text-to-3D optimization. We restrict to qualitative comparison for this task as quantitative metrics have high variance due to the absence of a ground truth dataset.
@@
"a hamburger" "a tulip" "a statue"
SDS
Refer to caption Refer to caption Refer to caption
SDI
Refer to caption Refer to caption Refer to caption
Ours
Refer to caption Refer to caption Refer to caption
@@ 5 Conclusion
We have demonstrated that compared to SDS, this method achieves better mode alignment as well as lower gradient variance, which in practice translates to more realistic optimization results as well as improved convergence rate.
Since this method simply provides optimization gradient much like SDS does, it may be used as a one-to-one replacement without retraining of the underlying model, or indeed substantial code modification.
While the basic algorithm works as the theory predicts in synthetic scenarios, with real-world models we have to contend with integrator error due to large score magnitudes. We have designed heuristics to alleviate this and achieve improvements on SDS in practice, but we hope future work will be able to improve the integration and/or sampling procedure, obviating the need for heuristics.
Impact Statement
@@ Appendix A Implementation details
       for $t=T-1,\ldots,0$ do
             $z_{t}\leftarrow\epsilon_{\theta}(z_{t},t,c)-(x-z_{t})/\lambda^{2}$
       end for
@@
       for $t=T-1,\ldots,0$ do
             $z_{T}\leftarrow z^{*}_{t}+(\epsilon_{\theta}(z_{t},t,c)-z^{*}_{t})e^{-\Delta t/\lambda^{2}}$
       end for
@@
Figure 7: Pseudocode of SDS and our procedure, MSD. We additionally show the numerically stable solver, which is used for experiments with Stable Diffusion.
Appendix B List of prompts
@@
Appendix C Ablations and more results
Figure 8: We extend Figure 2 with two ablations; applying guidance in the entire denoising trajectory (row 1) and noise scaled sample in the kernel term (row 2) (§ 4.3). We also show additional results for our method, SDI and SDS (row 3,4,5 respectively).

### v2

@@ 4.3 Practical Setting
For large-scale image datasets, idealized denoiser is no longer tractable and we contend with a learned denoising function, and the associated machinery. This introduces numerical issues. Namely, the magnitude of the kernel term may grow to where the standard first or second order integrators can no longer manage it (Section 3.4); but conversely, so does the magnitude of the learned score when $z_{t}$ is far out of distribution, because the ideal denoiser (Section 4.1) uses the same equation as mean shift. Start of the optimization, it is likely in a high-dimensional space that $x$ will be out of distribution and we have to choose between the integration failing because the denoiser term has a high magnitude, or because the kernel term has a high magnitude.
Table 3: Efficiency ($\uparrow$) on 2D toy density datasets. Left: ideal denoiser / Right: learned denoiser.
Fractal Spiral Pinwheel
SDS -7.37 / -6.89 -8.48 / -7.57 -7.82 / -6.99
VSD -5.92 / -3.83 -6.85 / -4.45 -6.36 / -3.93
SDI 14.18/ 14.21 13.94 / 14.17 14.19 / 14.18
Ours 13.44 / 7.65 13.38 / 6.32 13.76 / 7.08
@@ 4.4 Pre-trained Stable Diffusion
We use the latent-space diffusion model, Stable Diffusion, as the diffusion prior for text-conditioned optimization of parameters of differentiable image generators. Specifically, we optimize parameters $\vartheta$ of generator $g$, a rendering function that maps $\vartheta$ to an image $\mathcal{I}$. The rendered image $\mathcal{I}$ is fed to the image encoder to get $x^{k}$, our latent at optimization step $k$, over which the gradient is computed. We define two settings where $\vartheta$ (1) represents an RGB image, and (2) represents a 3D volume. Specifically:
@@
We implement all our code in PyTorch, on a single NVIDIA A100 gpu. We use the Threestudio (Guo et al., 2023) framework for experiments involving pre-trained Stable Diffusion. We use AdamW optimizer with lr$=10^{-2}$. We set optimization steps to $400$ for text-to-2D and $10k$ for text-to-3D. We use a monotonically decreasing schedule for the bandwidth $\lambda$.
@@
Figure 4: FID vs optimization iterations for text-to-2D generation.
@@ 4.5 Evaluation
We use a subset of the prompts curated by (Poole et al., 2022; Hertz et al., 2023). We include all prompts in Appendix D.
@@
For toy density dataset (Section 4.2), we compute negative log-likelihood scores (NLL), generative precision and recall (Kynkäänniemi et al., 2019), and maximum mean discrepancy (MMD). For text-to-2D, we use images produced by DDIM to represent the ground truth distribution. To evaluate fidelity of the images, FID (Heusel et al., 2017) is computed for each baseline (SDS, VSD, SDI) and ours against this ground truth image set. We also compute CLIP scores (cli, ) to measure prompt-generation alignment.
@@
Table 4 reports results for FID and CLIP-based similarity, comparing our method with SDS, VSD, and SDI. We outperform all baselines in image fidelity and achieve faster convergence, as measure via fid vs iterations in Figure 4.
@@
Figure 2 (top row) and Figure 8 compares our method with SDS, VSD, and SDI on text-to-2D generation, qualitatively. We show the importance of the two heuristics (Section 4.3) to resolve numerical instabilities, absence of which can result in visual artifacts in Figure 9. SDS, as discussed, produces low-fidelity results while SDI’s inversion accumulates numerical errors during early stages of optimization. In Figure 6, we qualitatively compare results for text-to-3D optimization. We restrict to qualitative comparison for this task as quantitative metrics have high variance due to the absence of a ground truth dataset.
@@ 5 Conclusion
We have demonstrated that compared to SDS, this method achieves better mode alignment as well as lower gradient variance, which in practice translates to more realistic optimization results as well as improved convergence rate. Since this method simply provides optimization gradient much like SDS does, it may be used as a one-to-one replacement without retraining of the underlying model, or indeed substantial code modification.
While the basic algorithm works as the theory predicts in synthetic scenarios, with real-world models we have to contend with integrator error due to large score magnitudes. We have designed heuristics to alleviate this and achieve improvements on SDS in practice, but we hope future work will be able to improve the integration and/or sampling procedure, obviating the need for heuristics, in-addition to using adaptive bandwidth annealing strategies.
@@
Acknowledgments
The authors would like to thank Pradyumn Goyal, Dmitry Petrov, Ashish Singh, and Artem Lukoianov for their helpful discussions and insights. This project has received funding from the European Research Council (ERC) under the Horizon research and innovation programme (Grant agreement No. 101124742).
@@ Appendix A Implementation details
       for $t=T,\ldots,1$ do
             $z_{t-1}\leftarrow\epsilon_{\theta}(z_{t},t,c)-(x-z_{t})/\lambda^{2}$
@@
       for $t=T,\ldots,1$ do
             $z_{t-1}\leftarrow z^{*}_{t}+(\epsilon_{\theta}(z_{t},t,c)-z^{*}_{t})e^{-\Delta t/\lambda^{2}}$
@@
Figure 7: Pseudocode of SDS and our procedure, MSD. We additionally show the numerically stable solver, $\dots$, stable), which is used for experiments with Stable Diffusion. Note, there is stochasticity in the ODESolver.
Appendix B Discussions
Why mode-seeking?
The desirability of mode seeking varies between applications. When trying to directly sample images from the trained model, we wish to sample from the full variety of the distribution instead of getting only the mode— we want sampling to interpolate between mode-seeking and mode-covering. Methods like DDIM aim for this. On the other hand, when we are optimizing an image (or using the image as a proxy to optimize, e.g. NeRF parameters), any gradient-based optimization will converge to a set of sparse points - local extrema - where the gradients are zero (if it converges at all). This is the intended use-case for SDS, VSD, SDI, and our method, and in this case, it is not possible in general to have the optimization process converge to a distribution of points. Given that, the best we can guarantee is that the points the process converges to are aligned with the distribution. Mode-seeking is our proposed way of achieving that.
Compatibility with other guidance schemes.
In low-dimensional settings (eg, our toy experiments), our method can recover the modes and reconstruct the data distribution well without any guidance (See Figures 1 and 3). This is aided by the fact that the conditional score estimates parameterized as $\epsilon_{\theta}(z_{t},c)$ (predicted noise from the pre-trained network) is good by itself, without guidance i.e. $\tilde{\epsilon}_{\theta}(z_{t},c)$. Empirically, we observe that without guidance, ancestral sampling techniques like DDIM produce samples that lie on the data manifold, albeit with few outliers.
This is not the case in the high-dimensional setting with experiments on Stable Diffusion. Here $\epsilon_{\theta}(z_{t},c)$ samples are noticeably bad and are predominantly outliers. Currently, the best fix is to augment these noise estimates with guidance to produce $\tilde{\epsilon}_{\theta}(z_{t},c)$, the strategy prevalent in sampling algorithms. We inherit these practices when performing distillation.
Guidance mechanisms alternative to CFG (Ho & Salimans, 2021) have been proposed, like Autoguidance (Karras et al., 2024). As these methods pair well with DDIM (and other ancestral sampling techniques), we believe the benefits will extend to distillation-based methods like ours. Ultimately from the perspective of distillation, different guidance simply changes the shape of the output distribution but does not fundamentally change the mechanics of diffusion. While we show results using Autoguidance in Table 1, we used CFG in all the remaining experiments as it is more widely used, has hyperparameters (guidance scale) that have been more rigorously tested by the community, and was used in all our baselines (SDS, VSD, and SDI).
Appendix C Ablations and More Results
Figure 8: We also show additional results for our method (full), SDI, VSD, and SDS.
Figure 9: We extend Figure 2 with two ablations; applying guidance in the entire denoising trajectory (row 1) and noise scaled sample in the kernel term (row 2) (§ 4.3).
Appendix D List of prompts
@@
Appendix E Licenses
Here we provide the URL, citations and licenses of the open-sourced assets we use in this work.
Table 5: URL, citations and licenses of the open-sourced assets we use in this work.
URL Citation License
https://github.com/threestudio-project/threestudio [39] Apache License 2.0
https://github.com/Stability-AI/stablediffusion [39] MIT License
https://github.com/NVlabs/edm2 [12] CC BY-NC-SA 4.0